### v1

# Experience Retrieval-Augmentation with Electronic Health Records Enables Accurate Discharge QA

Abstract
To improve the reliability of Large Language Models (LLMs) in clinical applications, retrieval-augmented generation (RAG) is extensively applied to provide factual medical knowledge. However, beyond general medical knowledge from open-ended datasets, clinical case-based knowledge is also critical for effective medical reasoning, as it provides context grounded in real-world patient experiences. Motivated by this, we propose Experience Retrieval-Augmentation (ExpRAG) framework based on Electronic Health Record (EHR), aiming to offer the relevant context from other patients’ discharge reports. ExpRAG performs retrieval through a coarse-to-fine process, utilizing an EHR-based report ranker to efficiently identify similar patients, followed by an experience retriever to extract task-relevant content for enhanced medical reasoning. To evaluate ExpRAG, we introduce DischargeQA, a clinical QA dataset with 1,280 discharge-related questions across diagnosis, medication, and instruction tasks. Each problem is generated using EHR data to ensure realistic and challenging scenarios. Experimental results demonstrate that ExpRAG consistently outperforms a text-based ranker, achieving an average relative improvement of 5.2%, highlighting the importance of case-based knowledge for medical reasoning.
Experience Retrieval-Augmentation with Electronic Health Records Enables Accurate Discharge QA
Justice Ou1 thanks: Correspondence to zo6@illinois.edu thanks: The two first authors made equal contributions., Tinglin Huang2 22footnotemark: 2, Yilun Zhao2 , Ziyang Yu3, Peiqing Lu, Rex Ying2 University of Illinois Urbana-Champaign1, Yale University2 , University of Waterloo3
1 Introduction
Figure 1: An illustrative example of utilizing experience from relevant clinical cases to support medical decision.
Benefiting from pretraining on large-scale corpora, Large Language Models (LLMs) are capable of performing complex reasoning and have shown great promise in medical applications Zheng et al. (2024); Liu et al. (2024). One important application is inferring clinical conditions, including diagnosis and medication, which can be formulated as a question-answering (QA) task Singhal et al. (2025); Chen et al. (2023); Huang et al. (2024). However, LLM agents often suffer from hallucinations and a lack of domain-specific knowledge, which limits their reliability in real-world medical applications.
To address this, prior studies have resorted to retrieving factual knowledge from open-ended databases to provide context, such as the description of drugs from Wikipedia Xiong et al. (2024); Yang et al. (2025). Such external knowledge enables LLMs to access general medical facts, thereby improving response accuracy. However, introducing such general facts cannot effectively help LLMs solve real clinical cases, which often involve coexisting clinical conditions. For example, as shown in Figure 1, adjusting a patient’s warfarin dosage requires reasoning based on the specific clinical context, whereas conventional retrieval can only provide the standard dosage for warfarin, which is irrelevant in this case.
In light of this, we argue that, in addition to general factual concepts, clinical case-based knowledge is also crucial for effective medical reasoning. The intuition is that an experienced clinician often relies on past cases with similar conditions to guide diagnosis, treatment decisions, and discharge planning. In this study, we propose Experience Retrieval-Augmentation (ExpRAG) framework, leveraging a large-scale EHR database MIMIC-IV Johnson et al. (2023b) as its knowledge basis. Specifically, ExpRAG breaks down the retrieval process into two steps: (1) report ranking applies an EHR-based similarity measurement to identify patients with similar medical conditions, and (2) experience retrieval extracts problem-relevant content from these patients’ discharge reports, which serves as the case-based contextual knowledge for LLMs. The introduced EHR modality enables large-scale clinical experience retrieval, grounding the model’s reasoning in real-world clinical practices.
To evaluate ExpRAG’s capability in medical reasoning, we introduce DischargeQA  [^1] Uploading to https://physionet.org/ , please check updates on Github: https://github.com/jou2024/EXPRAG, a clinical dataset including 1,280 QA pairs dedicated to discharge-related problems. The dataset primarily includes three types of problems, simulating the discharge process from initial diagnosis to medication prescription and post-discharge instructions. For each problem, we follow the data structure of the discharge report and select the content preceding the question as the problem background to avoid label leakage. Additionally, we index the option candidates using EHR to generate contextually relevant options, ensuring a non-trivial and clinically meaningful challenge for the model.
We evaluate the performance of four different LLMs and compare ExpRAG with the text-based report ranker using DischargeQA. The results demonstrate the effectiveness of using EHR to retrieve relevant clinical experience, as it consistently improves the performance of LLM backbones and outperforms the text-based ranker with an average relative improvement of 5.2%. Overall, our main contributions are listed below:
  • We propose ExpRAG, an EHR-based experience retrieval-augmentation framework, shedding light on the potential of leveraging past clinical cases to enhance LLM performance in medical reasoning tasks.
  • We introduce DischargeQA, a medical QA dataset for discharge-related questions, designed to evaluate LLMs’ ability to simulate the clinical decision-making process during patient discharge with a more challenging setup.
  • Our results demonstrate the advantage of ExpRAG over the text-based ranker, highlighting the effectiveness of EHR in providing clinically meaningful context.
2 Related Work
Retrieval-Augmented Generation (RAG) RAG has become a key paradigm for overcoming the static knowledge limitations of LLMs by retrieving external information Gu et al. (2018); Petroni et al. (2019). Traditional RAG frameworks typically use dense retrieval methods to augment generative tasks Devlin et al. (2019); Xiong et al. (2021). While effective in general QA, these methods often lack domain specificity, which is critical in healthcare Lu et al. (2024). Recent advancements like ClinicalRAG Lu et al. (2024) and MIRAGE Xiong et al. (2024) address this by integrating structured EHR data and clinical notes for diagnosis and treatment planning. However, existing benchmarks primarily focus on isolated information retrieval Johnson et al. (2023a); Kweon et al. (2024), overlooking the complexities of reasoning over patient histories and similar cases. To bridge this gap, our work extends RAG by combining structured EHR data with discharge summaries, enabling experience-driven reasoning for more realistic and reliable medical QA.
Medical QA Benchmark EHRSQL (Lee et al., 2022) and DrugEHRQA (Bardhan et al., 2022) target structured data queries, with the former addressing SQL-based operations and the latter focusing on drug-related questions. EHRNoteQA (Kweon et al., 2024) and RadQA (Soni et al., 2022) leverage clinician-verified QA pairs from discharge summaries and radiology reports, while MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019) evaluate LLMs with questions from medical exams or PubMed articles. Discharge-summary-focused datasets like emrQA (Pampari et al., 2018) and CliniQG4QA (Yue et al., 2021) use discharge notes for QA tasks, and specialized datasets like RxWhyQA (Fan, 2019) and drug-reasoning QA (Moon et al., 2023) focus on specific question types like medication reasoning. Unlike these benchmarks, DischargeQA introduces an evaluation framework centered around the discharge process, simulating the clinical workflow from diagnosis inference to medication prescription and discharge instruction generation. Additionally, we leverage EHR to generate non-trivial, contextually relevant candidate options, providing a more challenging and realistic setup.
Figure 2: The overview of ExpRAG: Given a medical query and the patient’s background, ExpRAG first indexes similar patients based on diagnosis, medication, and procedure similarity from the EHR. A text retriever is then applied to the discharge reports of the top-ranked similar patients to extract clinically relevant content, which is subsequently fed into the LLM to generate the answer.
3 ExpRAG Framework
ExpRAG provides a comprehensive framework for retrieving relevant knowledge from the cohort, as shown in Figure 2. We first formulate the problem that ExpRAG aims to tackle and elaborate the two-step retrieval framework including an EHR-based report ranker.
3.1 Task Formulation
A cohort contains a set of discharge report $\mathcal{D}=\{D_{i}\}_{N}$ where $D_{i}=\{d_{j}\}_{M}$ denotes the $i$-th report and $d_{j}$ is $j$-th paragraph in $D_{i}$. Each report is a medical document that offers an overview of a patient’s hospitalization. The goal of ExpRAG is to extract relevant content from $\mathcal{D}$ that helps LLM to effectively answer a given medical query $q$ related to a specific patient $p$:
$d_{*}=f_{\textsc{ExpRAG}}(p,q,\mathcal{D})$ (1)
The queries studied in this work focus on providing professional medical guidance for patients, including diagnosis, medication, and discharge instructions, thereby simulating realistic and practical clinical scenarios, as discussed in Section 4.
Different from the conventional RAG focusing on extracting factual concepts from open-ended databases, ExpRAG aims to utilize contextually-relevant clinical practice, inspired by how doctors collect and apply experience from past clinical cases. These two approaches rely on different reasoning procedures and knowledge sources, making them complementary to each other.
3.2 Coarse-to-Fine Retrieval Framework
To retrieve information from the cohort, one naive solution is to concatenate all the reports into one document and apply a text retriever to extract relevant content, similar to the conventional RAG pipeline. However, a standard EHR cohort typically contains millions of hospital visits, making it impractical to exhaustively search over all reports.
To efficiently perform experience retrieval, ExpRAG applies a two-step framework which conduct the retrieval from a coarse to fine level:
Report Ranking Before addressing a specific medical query, an intuitive assumption is that only patients with similar clinical histories, e.g., similar diseases or medications, can potentially provide meaningful guidance. In light of this, ExpRAG first employs a report ranker to efficiently discard unrelated cases and narrow down the candidate pool using the patient information:
$\mathcal{D}^{\prime}=f_{\text{Ranker}}(p,\mathcal{D})$ (2)
where $\mathcal{D}^{\prime}=\{D_{i}\}_{N^{\prime}\ll N}$ is a small subset of the selected discharge summaries. The ranker module will be scalable and enable effective utilization of patient context. Specifically, we introduce EHR as a knowledge basis to facilitate the patient-level similarity measurement, as presented in Section 3.3.
Experience Retrieval Based on the selected candidate pool, a sophisticated text retriever is capable of providing more accurate and dedicated clinical experience searching:
$d_{*}=f_{\text{Retriever}}(q,\mathcal{D}^{\prime})$ (3)
Built on top of clinically relevant reports identified by the dedicated ranking approach, the retriever focuses on extracting content related to the medical query. We here apply existing text retrievers, such as BM25 and Contriever, during this phase.
3.3 EHR-Based Report Ranker $f_{\text{Ranker}}$
Electronic Health Record (EHR), as a structured data organization, typically consists of multiple tabular data, each recording specific medical information about patients. In this study, we focus on measuring the similarity between patients using the following three medical entities:
  • Diagnosis: Identified disease assigned to a patient, represented by ICD-10 code.
  • Medication: Prescribed drugs administered to a patient, recorded using NDC code.
  • Procedure: Medical intervention, operation, or clinical process performed on a patient, represented by ICD-10 code.
Quantify the similarity between patients based on these three dimensions offers a comprehensive criterion for identifying clinically relevant reports.
For a patient $p$, these three medical entities are represented as sets $E^{\text{Diag}}_{p}$, $E^{\text{Med}}_{p}$, and $E^{\text{Proc}}_{p}$, respectively. Given two patients $p$ and $p^{\prime}$, we first compute the set similarity between them using each medical information:
$\tau_{\text{Diag}}=f_{\text{similarity}}(E^{\text{Diag}}_{p},E^{\text{Diag}}_{p^{\prime}}),$ (4)
$\tau_{\text{Med}}=f_{\text{similarity}}(E^{\text{Med}}_{p},E^{\text{Med}}_{p^{\prime}}),$ (5)
$\tau_{\text{Proc}}=f_{\text{similarity}}(E^{\text{Proc}}_{p},E^{\text{Proc}}_{p^{\prime}})$ (6)
where $f_{\text{similarity}}(\cdot,\cdot)$ is a set similarity metric, with the Jaccard Index applied in this study. Finally, these similarity metrics are aggregated using a weighted sum:
$\tau=\lambda_{1}\tau_{\text{Diag}}+\lambda_{2}\tau_{\text{Med}}+\lambda_{3}\tau_{\text{Proc}}$ (7)
where $\lambda_{1/2/3}$ is the hyperparameter balancing the importance of each metric. We perform pairwise similarity comparisons between the query patient and other patients within the EHR, returning the discharge summaries of the top-$k$ most similar patients as results.
Efficiency Analysis The overall computation is practically efficient since the computation of Jaccard Index can be significantly accelerated with some libraries, such as Faiss Douze et al. (2024) and NumPy Harris et al. (2020). Besides, indexing medical entities from tabular data enables fast lookups, further reducing computational overhead.
Task Answer Format #Query Example Question Practical Significance Background Source Option Source
Diagnosis Inference Multi-select 436 "Which diagnoses should be documented in the patient’s discharge summary?" Reflects a doctor’s process of identifying all relevant diagnoses based on clinical profile. Clinical profile Discharge Report & EHR
DischargeQA Medications Inference Multi-select 444 "Which medications should be prescribed to the patient at discharge?" Simulates the doctor’s task of ensuring correct medications are prescribed based on hospital treatment. Clinical profile & In-hospital progress Discharge Report & EHR
Instructions Inference Single-select 400 "What is the best instruction for this patient?" Mimics the final step of doctors’ advising patients with appropriate post-discharge care instructions. Clinical profile & In-hospital progress Discharge Report & AI
EHRNoteQA Clinical Inference Single-select & Open-Ended 962 "What was the treatment provided for the patient’s left breast cellulitis?" Extract and answer based on content from full discharge notes Full clinical notes Discharge Report & AI
CliniQG4QA Retrieval Text Span 1,287 "Why has the patient been prescribed hctz?" Retrieve the related content as answer from report Full clinical notes /
Table 1: Comparison of DischargeQA and the previous EHR-related QA benchmarks.
4 DischargeQA Dataset
To evaluate LLMs’ capability in utilizing the retrieved experience, we construct a medical question-answering dataset specifically designed for discharge-related queries based on MIMIC-IV Johnson et al. (2023b). Each question in the dataset pertains to critical discharge information, including the patient’s diagnosis, prescribed medications, and post-discharge instructions.
4.1 Dataset Introduction
Overview As shown in Table 1, DischargeQA consists of a total of 1,280 QA pairs, each associated with a patient ID and corresponding clinical background. The questions in DischargeQA can be categorized into three main types:
  • Diagnosis Inference: Questions related to identifying the patient’s medical diagnosis.
  • Medication Inference: Questions regarding the medications prescribed, including dosage, frequency, and purpose.
  • Instruction Inference: Questions focused on discharge instructions, such as follow-up care, activity restrictions, and self-care guidelines.
These three categories collectively cover the key aspects of discharge-related patient care, requiring LLMs to perform non-trivial reasoning based on the given clinical background. Moreover, all these problems can potentially benefit from the retrieved experience, mimicking the way clinicians apply past clinical knowledge to make medical decisions during discharge.
Figure 3: Inference pipeline of DischargeQA.
Evaluation Settings Each problem in DischargeQA includes a problem description, the patient’s clinical background, and multiple options for the LLM to choose from. While instruction inference uses a single-select setup, diagnosis and medication inference adopt a multi-select setup, requiring the model to select multiple options to answer the questions. Each option corresponds to a specific diagnosis or treatment. This multi-select format presents a more challenging and realistic setting for LLMs, as clinicians often need to identify and address multiple coexisting conditions.
The overall inference pipeline is presented in Figure 3, where we implement several components, including Ranker, Retrieval, and LLM agent, to support discharge-related QA using ExpRAG.
Comparison
Compared with existing QA benchmarks that focus primarily on general clinical QA or information retrieval, DischargeQA centers on the discharge procedure, simulating a doctor’s medical reasoning process: inferring the diagnosis from the clinical profile, prescribing appropriate treatments, and summarizing the condition—offering a more realistic scenario. Additionally, we utilize EHR to generate contextually relevant options, requiring LLMs to perform non-trivial reasoning to solve the tasks. More details are provided in Table 1.
4.2 Dataset Construction
Patients Filtering We first filter out low-quality patient records in MIMIC-IV for various reasons. Starting with all 430,000 patients, we remove encounters without available discharge summaries, leaving 320,000 patients. Next, we filter out patients with fewer than 3 or more than 40 entries in any of the diagnosis, medication, or procedure records, resulting in a final dataset of 28,000 patients for generating QA pairs. For instruction inference, we further exclude patients with excessively short discharge summaries using GPT-4o.
Background Generation To enable LLMs to make realistic medical decision, it is necessary to offer a clinical background of the patient along with the question. To avoid label leakage during the context generation, we propose leveraging the structured format of the discharge summary, which consists of the following components:
  • Clinical profile: Essential patient demography, the presenting condition, and initial clinical assessments.
  • In-hospital progress: The interventions, therapies, and the patient’s clinical progress during hospitalization.
  • Discharge plan summary: The details of discharge diagnosis and medication, and instructions for post-hospital care.
An illustrative example can be found in Appendix A.1. These three components represent the core of discharge decision-making procedure, aligning with the three main problem types in DischargeQA. Accordingly, we present the summarized sections preceding the questions as context. For example, the clinical profile serves as the background for diagnosis-related questions, while the in-hospital progress is additionally included for medication-related questions. Notably, basic patient demographic information is always included as part of the contextual background.
Option Generation For all tasks, the correct options are directly extracted from the discharge reports. For diagnosis and medication inference, we utilize EHR to generate non-trivial candidate options by extracting all associated diagnoses and medications of a patient and feeding them into GPT-4o to select contextually relevant candidates, ensuring a challenging selection process. For instruction inference, GPT-4o first summarizes the key points of the ground-truth answers and then applies permutations to generate plausible yet incorrect candidate options.
5 Experiments
In this section, we evaluate ExpRAG on DischargeQA with four LLMs, comparing it with the text-based report ranker. We also analyze the effects of balancing coefficients, the number of similar patients, and retriever choices, followed by case studies on the retrieved experience.
Model Report Ranker Instruction Diagnosis Medication
Acc(%) Acc(%) F1 Acc(%) F1
GPT-3.5 Direct-Ask 73.0 18.50 0.498 1.15 0.234
Text-based 78.8 15.60 0.405 0.68 0.317
EHR-based 79.5 18.81 0.504 1.80 0.371
GPT-4o Direct-Ask 90.0 9.86 0.510 3.65 0.486
Text-based 90.3 8.26 0.493 4.95 0.601
EHR-based 91.3 21.33 0.530 9.68 0.638
Deepseek-R1 Direct-Ask 70.0 10.78 0.363 0.69 0.217
Text-based 72.0 12.84 0.381 2.03 0.230
EHR-based 75.3 11.01 0.379 1.13 0.241
Mistral-7b Direct-Ask 67.0 16.03 0.488 1.13 0.362
Text-based 70.0 14.67 0.490 0.90 0.344
EHR-based 69.0 13.79 0.505 1.13 0.353
Table 2: Performance Comparison Across Multiple LLMs using DischargeQA
5.1 Comparison of LLMs
We evaluated the performance of 4 state-of-the-art LLMs of varying scales, ranging from close-source to 8 billion parameters open-source model, on three clinical tasks: discharge instructions, diagnosis, and medications. These models included (1) Proprietary LLMs GPT-3.5 OpenAI (2022) and GPT-4o OpenAI (2024)) (2) Open-source LLMs Mistral Jiang et al. (2023a), and Deepseek-R1 Guo et al. (2025).
Hyperparameters We also present the results of LLMs with ExpRAG. The default balancing coefficients $\lambda_{1/2/3}$ are set to 1/3 each and auto-merging Liu (2022) is used as the retriever. The number of similar patients is set as 15.
Metrics We report accuracy across all the tasks, calculated as the percentage of correctly answered questions. Note that for multi-select problems, a question is considered answered correctly only when all correct options are selected. We additionally report F1 scores for the multi-select problems to provide a more comprehensive analysis of LLM performance on these challenging tasks.
Results Table 2 summarizes the performance of all LLMs with ExpRAG on the three clinical tasks. GPT-4o consistently achieves the best performance across tasks, with an absolute improvement of 13% over GPT-3.5 on instruction inference problems. Additionally, the multi-select tasks (diagnosis and medication) prove significantly more challenging, as most LLMs achieve accuracy below 20%, indicating the limited medical reasoning capabilities of current large language models.
5.2 Comparison of Report Ranker
To verify the effectiveness of ExpRAG and the utilization of EHR, we introduce a baseline that performs report ranking solely based on text, i.e., a text-based ranker. We embed all text components, including the question, options, and background, into a single representation. Similarly, each discharge report is embedded, and similarity is computed with the question embedding. The top-$k$ reports are then retrieved, followed by a text retriever to extract the relevant information.
The results are presented in Table 2. We observe that ExpRAG outperforms the text-based ranker in most cases, achieving an average relative improvement of 5.2%. Notably, the EHR-based ranker leverages structured EHR data for ranking, eliminating the need for an embedding process and thereby enabling a more efficient pipeline.
5.3 Further Analysis
We conducted additional studies to investigate the impact of key components on the performance of ExpRAG, including the number of similar patients $k$, the balancing coefficients $\lambda_{1/2/3}$, and the choice of retrievers. We apply GPT-3.5 as backbone by default.
Model Instruction Diagnosis Medication
Acc Acc F1 Acc F1
Uniform 79.5 18.81 0.504 1.8 0.371
Task-focused 77.5 10.91 0.377 0.91 0.322
Complementary 76.8 18.18 0.446 2.73 0.305
Table 3: Performance Comparison for Coefficients Distributions DischargeQA
Balacing Coefficients As shown in Equ. 4, we introduce three coefficients to balance the similarity computed based on diagnosis, medication, and procedures. We apply an equal distribution by default and explore the effect of using different weighting strategies here:
  • Task-focused weighting: Assign a weight of 1 to the task-relevant similarity measure and 0 to the others. For example, $\lambda_{1}=1,\lambda_{2}=0,\lambda_{3}=0$ for diagnosis inference.
  • Complementary weighting: Assign a weight of 1 to the two less relevant similarity measures while setting the task-relevant measure to 0. For example, $\lambda_{1}=0,\lambda_{2}=1,\lambda_{3}=1$ for diagnosis inference.
The results are shown in Table 3. We can find that complementary weighting can achieve the best performance in most cases, demonstrating that information from multiple clinical dimensions can provide a more comprehensive context.
Top-$k$ Patients We vary the number of retrieved similar patients $k$ on QA performance. As shown in Table 4, different tasks have varies trends. For example, on Instruction task, accuracy starts to increases with larger number, from $k=10$ with an accuracy of 78.75. While for Diagnosis and Medication tasks, beyond $k=20$, performance fluctuates, suggesting that while more retrieved candidates provide useful context, excessive retrieval may introduce irrelevant or conflicting information, leading to slight declines in accuracy.
# Similar Patients Instruction Diagnosis Medication
Acc Acc F1 Acc F1
$k=5$ 80.00 19.04 0.511 1.80 0.366
$k=10$ 78.75 18.58 0.511 1.35 0.371
$k=15$ 79.50 18.81 0.504 1.80 0.371
$k=20$ 80.75 19.50 0.524 1.80 0.377
$k=25$ 82.25 19.27 0.515 1.35 0.352
Table 4: Performance Comparison with Different $k$ for GPT3.5.
Experience Retriever We evaluate four retrievers for experience retrieval: auto-merging Liu (2022), sentence-window Liu (2022), BM25 Robertson et al. (2009), BM25+ (the combination of BM25 and word embeddings), Contriever Izacard et al. (2021), and flare Jiang et al. (2023b). More details regarding the methods can be found in Appendix A.2.
As shown in Table 5, auto-merging, sentence-window, and flare achieve the best performance, highlighting the effectiveness of context-aware retrieval. However, Contriever, which utilizes unsupervised dense representations, also underperforms in our cases, suggesting the need for medical domain-specific fine-tuning.
Retriever Instruction Acc
Auto Merging 79.5
Sentence Window 74.5
BM25 68.0
BM25+ 69.0
Flare 74.5
Contriever 69.0
Table 5: Performance Comparison of Retrievers.
5.4 Case Studies
We conduct a case study focusing on Discharge Diagnosis in DischargeQA. We examine the discharge diagnosis of a target patient (ID: 20453584) who presented with bilateral ulnar paresthesias and neck pain. We compare this patient with similar patients, who are selected by comparing ICD/NDC codes from EHR tabular data. Fig 4 shows one example question in DischargeQA and the similarities in the discharge reports between the target patient and the similar patients with IDs 25633130, 29378221 and 28817667.
Figure 4: Comparison of Similar Patients.
Analysis of Similar Patients’ Data: Upon reviewing the discharge summaries of the similar patients, it became clear that several key diagnostic features were shared with patient 20453584:
  • Disc Herniation: Both the target patient and similar patients had disc herniations, with the target patient experiencing a C5-6 disc-osteophyte complex and the similar patients exhibiting C3-C4 and C6-C7 herniations.
  • Spinal Stenosis: Many of the similar patients displayed spinal stenosis, which was consistent with the target patient’s symptoms of narrowing of the spinal canal and foraminal narrowing.
  • Upper Extremity Symptoms: The target patient reported bilateral ulnar paresthesias, which mirrored the bilateral symptoms observed in several similar patients, such as neck pain radiating to the arms and tingling in the extremities.
Results: By comparing the discharge summaries, key features from similar patients that influenced the diagnosis of the target patient:
  • Similar patients with C6-C7 disc herniation and radiculopathy helped to refine the target patient’s diagnosis, suggesting that similar nerve root involvement could explain the upper extremity symptoms.
  • The presence of spinal stenosis and neural foraminal narrowing in several patients guided the understanding of the target patient’s potential nerve compression, which contributed to the diagnosis of spinal stenosis.
The comparison to similar patients led to a more precise discharge diagnosis for the target patient, which included a C5-6 disc-osteophyte complex with associated spinal canal narrowing and neural foraminal narrowing. These insights allowed the LLMs to confirm the target patient’s diagnosis, which aligned with options A and Gdisc osteophyte and spinal stenosis.
6 Conclusion
Inspired by the importance of experience in clinical decision-making, we propose a novel coarse-to-fine retrieval framework, ExpRAG, to utilize knowledge from similar patient records. Specifically, we introduce EHR as a knowledge basis and employ a comprehensive similarity measurement algorithm to narrow down the candidate pool, from which a text retriever extracts the relevant content. Evaluated on our curated DischargeQA, ExpRAG consistently improves the performance of various LLMs, highlighting the potential of leveraging past clinical experience for enhanced medical QA.
Limitations While EHR provides abundant medical information, such as lab test results, our proposed ExpRAG currently utilizes only diagnosis, medication, and procedures as an initial exploration. Additionally, DischargeQA currently consists solely of multi-option questions, which may limit its ability to comprehensively evaluate the generative capabilities of LLMs.
Appendix A Appendix
A.1 Dataset Details
MIMIC-IV [^2] https://mimic.mit.edu/docs/iv/ This database Johnson et al. (2016, 2023b, 2023a) contains information on over 40,000 patients admitted to the critical care units at Beth Israel Deaconess Medical Center from 2001 to 2012 and has been widely used in prior research. The database is publicly available for research purposes, with strict de-identification protocols to protect patient privacy, making it a valuable resource for developing and evaluating machine learning models in healthcare. The data is hierarchically organized, with each patient record comprising multiple encounters, each containing various entities such as demographics, medications, diagnoses, procedures, and lab results. Additionally, the database includes unstructured data, such as discharge reports and X-ray images, with each admission marked by a date and timestamp.
Structure of Discharge Report As shown in Figure 5, a discharge summary follows a structured format that provides a comprehensive overview of a patient’s hospitalization and care journey. It is typically divided into three main parts: (1) The Clinical Profile, which includes essential patient information, presenting conditions, and the initial clinical assessment upon admission; (2) The In-Hospital Progress, which documents the treatment plan, administered therapies, and the patient’s progress throughout the hospital stay; and (3) The Discharge Plan Summary, which summarizes the patient’s discharge status, prescribed medications, and detailed post-discharge instructions for ongoing care.
Figure 5: An example of discharge report, which can be split into 3 sections and 7 subsections: clinical profile, in-hospital progress, and discharge plan summary. Note: Some pertinent results from exams are before diagnosis, while some are after diagnosis or after procedures.
A.2 Retriever
Auto-merging Auto-merging retrieval in RAG by LlamaIndex Liu (2022) hierarchically structures documents into parent and child nodes, allowing for the retrieval of larger, more coherent context by merging child nodes into parent nodes when multiple related chunks are relevant to a query
Sentence-window Sentence-window retrieval by LlamaIndex Liu (2022) parses documents into individual sentences with surrounding context, enabling fine-grained retrieval while maintaining local coherence by including a window of adjacent sentences
BM25 & BM25+ BM25 (Best Match 25) is a ranking function used in information retrieval that scores documents based on the frequency of query terms within them, taking into account document length and term frequency saturation. BM25+ by LlamaIndex Liu (2022) combines retrieval methods BM25 and vector-based retrieval, to leverage the strengths of both approaches. This hybrid technique allows for capturing both keyword relevance and semantic similarity, often using algorithms like Relative Score Fusion to re-rank and merge results from different retrievers, resulting in more accurate and comprehensive search outcomes.
flare FLARE (Forward-Looking Active REtrieval) enhances RAG by enabling the language model to anticipate future content needs, iteratively predicting upcoming sentences and retrieving relevant information when encountering low-confidence tokens, thus improving response accuracy and contextual relevance
Contriever Contriever is a single-tower dense retrieval model that employs self-supervised contrastive learning to enhance document embeddings for retrieval tasks. It encodes both queries and documents using the same encoder, producing dense vector representations. The model utilizes a self-supervised contrastive learning approach with a loss function that optimizes embeddings by comparing relevant passages to negative (irrelevant) ones
A.3 LLM Backbone
GPT3.5 & GPT-4o GPT-3.5, developed by OpenAI and released in November 2022, was followed by GPT-4o, also created by OpenAI and launched on May 13, 2024, marking a significant advancement with its ability to process and generate outputs across text, audio, and image modalities in real time.
Deepseek-R1 DeepSeek-R1 is a powerful 671 billion parameter language model developed by DeepSeek AI, from which we use DeepSeek-R1-Distill-Llama-8B was derived as a more efficient 8 billion parameter version, distilled from the original model’s knowledge to offer improved performance on reasoning tasks while maintaining computational efficiency
Mistral Mistral-7B-Instruct-v0.3 is an advanced instruction-tuned language model featuring an extended vocabulary of 32,768 tokens, support for the v3 Tokenizer, and function calling capabilities, enabling more versatile and complex interactions compared to its predecessors
A.4 Prompt Templates
Figure 6 presents the prompt used to generate Diagnosis task options, which is the same way as generating options for Medication task.
Figure 6: Prompt design example for Diagnosis tasks